OpenStudy (anonymous):

Which pair of numbers is relatively prime? A. 10 and 33 B. 14 and 63 C. 26 and 91 D. 18 and 45

OpenStudy (anonymous):

@Qwertty123

zepdrix (zepdrix):

Ooo they gave you some tricky options :P Not even a pair of even numbers lol hmm

OpenStudy (anonymous):

i honestly have no idea ;-;

zepdrix (zepdrix):

Try to break the numbers down a little bit :) 14 = 7*2 63 = ?*?

OpenStudy (anonymous):

idek what that means XD

zepdrix (zepdrix):

We would like to rewrite each number as multiplication. And see if any of the multiplication numbers match up :)

zepdrix (zepdrix):

Example: 6 = 3*2 8 = 4*2 6 and 8 are NOT relatively prime, because they both have a 2 as a multiplier

zepdrix (zepdrix):

14 = 7*2 63 = ?*? what two numbers give you 63? :o

OpenStudy (anonymous):

3 and 31 o-o ?

zepdrix (zepdrix):

Hmm 3*31 = 93 woops :O

OpenStudy (janu16):

9 and 7

OpenStudy (anonymous):

3 and 23

zepdrix (zepdrix):

There is a nice divisibility trick for 3's and 9's. If the digits add up to a multiple of 9, then it's divisible by 9. 6+3 = 9, so it's 9*something and ya you spilled the beans -_- le sigh..

zepdrix (zepdrix):

It would've been 3*21. But then we have to break down the 21 further D: So anyway, 14 = 7*2 63 = 7*9 They both have 7 as a factor/multiplier, so they are NOT relatively prime. So b is NOT the option we're looking for.

zepdrix (zepdrix):

How bout option A? 10 = ?*? 33 = ?*?

OpenStudy (anonymous):

isnt it c?

zepdrix (zepdrix):

26 = 13*2 91 = 13*7 91 is a kind of difficult to break down, but it turns out they both have a 13 in them.

OpenStudy (anonymous):

so its right?

zepdrix (zepdrix):

if they share something... then they are NOT relatively prime. So no, it's not right

OpenStudy (anonymous):

>-> crap

zepdrix (zepdrix):

Try option A maybe, Break down the 10 into factors, and also the 33

OpenStudy (anonymous):

5*2 and 3*11

OpenStudy (anonymous):

?

zepdrix (zepdrix):

Ok good :) 5*2 3*11 Oooo I think we did it!! ya? The second set of numbers has no 5 or 2 in it, and the first set of numbers has no 3 or 11 in it. So they are relatively prime! :)

OpenStudy (anonymous):

thanks so much